A father is now twice as old as his daughter. 10 years ago he was 3 times as old as her. How old are they both now?
Answer by J2R2R(94)

f is age of father; d is age of daughter
Now, f = 2d
Ten years ago
(f - 10) = 3(d - 10)
Replacing f with 2d and expanding
2d - 10 = 3d - 30
20 = d
f = 40
So the father is now twice the age of his daughter (40 = 2 x 20)
Ten years ago the father was three times the age of his daughter (30 = 3 x 10)
The answer is the father is now 40 years and the daughter is 20 years.